OpenStudy (anonymous):

What is the inverse of y=9x^2 -4

OpenStudy (kittiwitti1):

http://mathinsight.org/definition/inverse_function

OpenStudy (anonymous):

First what you want to do replace every x with a y and replace every y with an x so first you have y=9x^2-4 then you'll have x=9y^2-4 now get y to be alone to get y alone, first add 4 to both sides so you get x+4 = 9y^2 now you can divide both sides by 9, so y can be alone so you'll have (x+4)/(9) = y^2 now to get rid of the square from the y, take the square root of both sides, so you'll have sqrt(y^2) = sqrt((x+4)/(9)) the square root cancels out the power 2 in the y so you get y = sqrt((x+4)/(9)) which can equal to y=sqrt((x+4))/(3) (since we did the square root of 9 is 3, you dont put a square root over the denominator anymore)

OpenStudy (marcelie):